OpenStudy (anonymous):

double integral 3x^2-3y^2 dydx on circle x^2+y^2=4 how do i find the endpoints using the circle.

OpenStudy (anonymous):

convert to polar coordinates. \[x=r*\cos(\theta); y=r*sin(\theta)\] \[F(x,y)=F(r,\theta)=3r^2cos^2(\theta)-3r^2sin^2(\theta)\] your limits of integration would be \[r=[0,2]; \theta=[0.2\pi]\] the integral would then become \[\int\limits_{0}^{2\pi}\int\limits_{0}^{2}r*F(r,\theta)drd\theta\]

OpenStudy (nubeer):

end points of r?

OpenStudy (nubeer):

i meant the limit?